OpenStudy (anonymous):

Use the sum-to-product formulas to write the sum or difference as a product. sin 5θ + sin 3θ

OpenStudy (ranga):

sin(A) + sin(B) = 2sin((A+B)/2) * cos((A-B)/2) Here A = 50, B = 30

OpenStudy (anonymous):

I don't understand.

OpenStudy (anonymous):

Sin5x+Sin3x = 2sin((5x+3x)/2) cos((5x-3x)/2 = 2siin8x/2 cos2x/2 = 2sin4x cosx